### v1

# IRONIC: Coherence-Aware Reasoning Chains for Multi-Modal Sarcasm Detection

Abstract
Interpreting figurative language such as sarcasm across multi-modal inputs presents unique challenges, often requiring task-specific fine-tuning and extensive reasoning steps. However, current Chain-of-Thought approaches do not efficiently leverage the same cognitive processes that enable humans to identify sarcasm. We present IRONIC, an in-context learning framework that leverages Multi-modal Coherence Relations to analyze referential, analogical and pragmatic image-text linkages. Our experiments show that IRONIC achieves state-of-the-art performance on zero-shot Multi-modal Sarcasm Detection across different baselines. This demonstrates the need for incorporating linguistic and cognitive insights into the design of multi-modal reasoning strategies. Our code is available at: https://github.com/aashish2000/IRONIC
IRONIC: Coherence-Aware Reasoning Chains for
Multi-Modal Sarcasm Detection
Aashish Anantha Ramakrishnan1thanks: These authors contributed equally to this work., Aadarsh Anantha Ramakrishnan211footnotemark: 1, Dongwon Lee1 1The Pennsylvania State University; 2National Institute of Technology, Tiruchirappalli 1{aza6352, dul13}@psu.edu, 2106121001@nitt.edu
1 Introduction
Figure 1: Overview of the IRONIC proposed framework
Sarcasm can be defined as a form of verbal irony where the goal is to mock or convey contempt while the literal meaning of the words is different, often opposite to the intended meaning Camp (2012). Multi-modal sarcasm utilizes the interplay between these visual and text cues to convey these intents. The task of multi-modal sarcasm detection is helpful for various applications, including social media analysis, sentiment analysis, and human-computer interaction Pang and Lee (2008). With traditional supervised learning approaches demonstrating limited success in this domain, recent research has focused on leveraging Multi-modal Large Language Models (MLLMs) to improve performance Tang et al. (2024). These models have shown promise even in Zero-Shot settings, where they can generalize to new tasks without explicit training data. The complexity of multi-modal sarcasm detection stems from the fact that it incorporates a combination of semantic, pragmatic, and analogical cues, which are not aligned with the logical and sequential nature of traditional reasoning strategies of language models Yao et al. (2025); Kumar et al. (2022). This leads to models taking shortcuts such as over-reliance on uni-modal cues and failing to discern sarcasm from other forms of figurative language Liang et al. (2022); Qin et al. (2023).
Existing approaches for prompt-based reasoning in MLLMs have focused on leveraging different varieties of logical chains, such as Chain-of-Thought (CoT) prompting Wei et al. (2022). Inspired by the success of text-only cognition-aware and linguistics-based reasoning strategies Wang and Zhao (2024); Lee et al. (2025), we propose IRONIC (explIcit Reasoning with cOherence relatioNs as cognItive struCtures). IRONIC leverages multi-modal Coherence Relations Anantha Ramakrishnan et al. (2025a); Alikhani et al. (2020) to provide a structured reasoning pathway for MLLMs to follow while identifying image-text incongruencies. Coherence Relations (CR) are a formal theory of discourse Hobbs (1978) that provides a formal framework for understanding how different communicative components relate to each other Hobbs (1978). As a task-independent framework, CR-based reasoning can help generate informative rationales for multi-modal sarcasm detection.
We analyze the capabilities of both proprietary (GPT-4o) and open-source models (InternVL3) in using Coherence Relations for reasoning Zhu et al. (2025); OpenAI et al. (2024). Our findings on two different datasets: MMSD2.0 Qin et al. (2023) and RedEval Tang et al. (2024) show that the higher-order reasoning abilities of MLLMs play a crucial role in how they leverage CRs. IRONIC enables more pragmatically-aware models such as GPT-4o OpenAI et al. (2024) to achieve state-of-the-art performance on zero-shot multi-modal sarcasm detection tasks. The contributions of this work can be summarized as follows:
  • We present IRONIC, a coherence-aware in-context reasoning framework that leverages Multi-modal Discourse Coherence Relations for effectively analyzing image-text incongruencies.
  • IRONIC enables MLLMs with emergent higher-order reasoning capabilities to outperform existing multi-step reasoning strategies such as CoT in zero-shot settings across different datasets.
  • Our results demonstrate the importance of leveraging linguistic and cognitive insights into the design of multi-modal reasoning chains, improving the performance of MLLMs on pragmatic and analogical comprehension tasks.
2 Related Work
2.1 Multi-Modal Sarcasm Detection
Initial approaches for multi-modal sarcasm detection utilized supervised deep learning models to incorporate image and text features Schifanella et al. (2016); Das and Clark (2018). Attention mechanisms further improved their abilities to associate image-text features across modalities Pan et al. (2020); Wu et al. (2021). This led to the incorporation of more modalities such as audio, video, and facial expressions Castro et al. (2019); Ray et al. (2022). However, models still struggled to comprehend complex relationships between image-text modalities required for sarcasm understanding. With the popularity of pre-training architectures such as CLIP Radford et al. (2021) and MLLMs, the capabilities of these models in identifying sarcastic content expanded further Jia et al. (2024). Prompting-based approaches gained prominence due to their ability to generalize across domains without task-specific fine-tuning Wang et al. (2024); Zhang et al. (2025b).
Relations Referential Pragmatic Analogical
Insertion
Concretization
Projection
Restatement
Extension
Table 1: Classification of CRs by type of cognitive reasoning. Green checkmarks indicate the presence of each reasoning type, while dots indicate absence.
2.2 Image-Text Relation Understanding
Traditional multi-modal models worked solely on the assumption that information across modalities is highly correlated with each other Radford et al. (2021), with tasks such as image captioning and image generation leveraging their strong semantic overlap Mahajan and Roth (2020); Liu et al. (2021). However, as these multi-modal models expanded their applications across different task domains such as news Anantha Ramakrishnan et al. (2024) and social media Xu et al. (2022); Vempala and Preoţiuc-Pietro (2019), the need for enhanced reasoning beyond semantic overlap became clear. Coherence Relations offer a cohesive framework for connecting different types of reasoning across multiple modalities and have improved model comprehension capabilities across different domains Alikhani et al. (2020, 2022); Xu et al. (2022); Sosea et al. (2021). Motivated by these insights, we analyze the incorporation of CRs as a cognitive reasoning scaffold for multi-modal sarcasm detection.
3 Methodology
The proposed IRONIC framework, as shown in Figure 1, constructs a reasoning chain with 2 main components: (1) Relation Extraction (2) Interpretative Reasoning. The predicted CR distributions and specific prompts are shown in the Appendix Sections A & C.
3.1 Relation Extraction
The Relation Extraction step analyzes each image-text pair and predicts the most applicable CR along with a rationale to justify its selection. As depicted in Table 1, we utilize CRs that provide coverage of referential, analogical, and pragmatic reasoning types with varying levels of granularity. The selection of these CRs is supported by their applicability to analyzing image-text pairs from social media domains Xu et al. (2022), ensuring strong generalizability with the datasets utilized for multi-modal sarcasm detection. The outputs of this step include the predicted CRs and their supporting rationales for each sample. Our framework incorporates 5 distinct CRs:
  • Insertion: The reader must infer the referent object or entity from the visual concept with no explicit mentions in the text.
  • Concretization: Identifies explicit referential alignment between image and text along with implied communicative intents.
  • Projection: Presents an analogical mapping between topics discussed in the text and the objects presented in the image.
  • Restatement: Describes an explicit referential agreement with strong semantic overlap across modalities.
  • Extension: Introduces narrative or storytelling metaphors in the text that convey speaker intent and elaborate further on the visual content.
3.2 Interpretative Reasoning
Although MLLMs show limited ability in explicit coherence understanding tasks such as predicting CRs Anantha Ramakrishnan et al. (2025a); Jia et al. (2025), recent studies have shown that language models are effective in utilizing coherence-aware reasoning chains for downstream higher-order reasoning tasks Zhang et al. (2025a); Anantha Ramakrishnan et al. (2025b). This motivates the second step of our reasoning chain, where we integrate the image-text pair, the predicted CR, and the generated rationale for distinguishing if a sample contains multi-modal sarcasm. In addition to these inputs, we provide task-specific instructions to discern simple humor or satirical commentary from sarcasm indicators such as mockery or irony concerning a specific referent. This enables our Interpretative Reasoning step to be robust, where the coherence-aware rationale is used as a supporting element while basing the final decision on the full multi-modal input. Thus, by providing both general coherence rationales and task-specific instructions, we develop a flexible prompting strategy that can handle inconsistencies in upstream reasoning steps.
Dataset Model Type Accuracy ↑ Precision ↑ Recall ↓ F1 ↑
MMSD 2.0 GPT4o Zero-Shot 0.7107 0.7952 0.7107 0.7024
Zero-Shot CoT 0.7468 0.8096 0.7468 0.7430
$S^{3}$ CoT 0.6559 0.7863 0.6559 0.6344
IRONIC 0.7659 0.7729 0.7659 0.7670
InternVL3-14B Zero-Shot 0.7393 0.7427 0.7393 0.7322
Zero-Shot CoT 0.7406 0.7752 0.7406 0.7400
$S^{3}$ CoT 0.6874 0.7727 0.6874 0.6775
IRONIC 0.7177 0.7294 0.7177 0.7032
RedEval GPT4o Zero-Shot 0.7799 0.8423 0.7799 0.7807
Zero-Shot CoT 0.7669 0.8290 0.7669 0.7677
$S^{3}$ CoT 0.7002 0.8172 0.7002 0.6939
IRONIC 0.8406 0.8501 0.8406 0.8422
InternVL3-14B Zero-Shot 0.8327 0.8319 0.8327 0.8309
Zero-Shot CoT 0.7420 0.8039 0.7420 0.7428
$S^{3}$ CoT 0.7122 0.8201 0.7122 0.7075
IRONIC 0.8197 0.8188 0.8197 0.8175
Table 2: Results for Multi-Modal Sarcasm Detection on both datasets
4 Experiments
4.1 Datasets
We select two popular multi-modal sarcasm detection datasets: MMSD2.0 Qin et al. (2023) and RedEval Tong et al. (2024) for evaluating the effectiveness of IRONIC. It builds on the popular Twitter-based MMSD dataset Cai et al. (2019) by removing spurious cues such as hashtags and unreasonable labels through re-annotation. MMSD2.0’s test set contains 1037 and 1372 examples for the sarcasm and the non-sarcasm classes. RedEval was constructed for Out-of-Domain evaluation of models trained on Twitter-based datasets. Samples in RedEval are sourced from Reddit, with 395 sarcastic examples selected from the "sarcasm" subreddit and 609 negative examples from other subreddits.
4.2 Evaluation Metrics
To measure the performance of MLLMs on multi-modal sarcasm detection, we use the F1 score (weighted) as our main metric since it is framed as a binary classification task. We also report overall Accuracy, Precision, and Recall for all models.
4.3 Baselines Compared
To understand how MLLMs utilize different types of reasoning chains compared to IRONIC, we include a variety of prompting strategies in our experiments. Zero-Shot serves as a simple baseline where only the input image and text are provided with a single task prompt for sarcasm prediction. Zero-Shot CoT requires the model to generate a step-by-step rationale before deciding whether sarcasm is present in the inputs. We also compare $S^{3}$ CoT, a strategy based on $S^{3}$ Agent Wang et al. (2024) for multi-modal sarcasm detection. $S^{3}$ Agent analyzes image-text pairs from 3 critical perspectives: superficial expression, semantic information, and sentiment expression using a multi-agent framework. To ensure consistency, we combine the parallel agent prompts that generate 3 separate rationales into one single prompt, followed by a decision agent to predict the presence of sarcasm. This way, $S^{3}$ CoT is also a two-step strategy, serving as an alternate reasoning chain baseline.
5 Results
IRONIC Improves over CoT Baselines
From our presented results on both datasets, we observe that IRONIC outperforms $S^{3}$ CoT across all settings and Zero-Shot CoT in 3 out of the 4 settings compared. With GPT-4o, we see an improvement of 3.23% and 9.70% on both MMSD2.0 and RedEval, respectively, compared to Zero-Shot CoT. Similarly, for InternVL3, we observe a 10.06% improvement over Zero-Shot CoT in RedEval. IRONIC when combined with GPT-4o beats all previous state-of-the-art Zero-Shot approaches for MLLM-based sarcasm detection on both datasets Zhang et al. (2025b); Tang et al. (2024). This demonstrates the benefits of a structured cognitive framework for multi-modal sarcasm prediction over empirical prompting strategies in Zero-Shot settings, supporting our initial hypothesis.
Challenges in Explicit Reasoning Incorporation
An interesting observation that we find when testing InternVL3-14B, is the strong zero-shot performance without separate reasoning steps on both datasets. The introduction of reasoning chains, both general and coherence-aware, does not provide meaningful performance improvements. This suggests that the model may rely on implicit heuristics for sarcasm recognition. Another potential reason for this behavior may be the lack of capacity to generate high-quality intermediate rationales Anantha Ramakrishnan et al. (2025a). This indicates that larger proprietary models may demonstrate higher internal reasoning fidelity to make use of cognitive reasoning scaffolds such as CRs.
6 Conclusion
We propose IRONIC, a coherence-aware in-context prompting strategy inspired by the Theory of Discourse Coherence for multi-modal sarcasm detection. Our study covers both proprietary and open-source models across image-text pairs sourced from 2 different social media platforms. We observe that IRONIC elicits improved reasoning capabilities from MLLMs compared to other CoT approaches. Our experiments also investigate the distinction between implicit pattern matching and explicit reasoning, demonstrating the importance of expanding model capabilities to include pragmatic and analogical reasoning, moving beyond factual and logical approaches.
Limitations
Our current experimental setup for IRONIC is limited to a couple of top-performing proprietary and open-source models. We are working on expanding our study to cover a more comprehensive set of models as a part of future work. Additionally, our evaluation strategy does not incorporate task-specific fine-tuning. Exploring multi-task fine-tuning strategies to improve both CR prediction and sarcasm recognition are promising research directions we aim to pursue. Finally, to facilitate a deeper analysis on the Relation Extraction step of IRONIC, we plan to explore human evaluation approaches for rationale quality analysis, facilitating identification of reasoning gaps and hallucinations.
Ethics Statement
With Multi-modal Sarcasm detection requiring implicit assumptions that leverage a model’s world view, we acknowledge the potential negative impacts of unintended stereotypical biases learned by these models. Similarly, the types of CRs we analyze are optimized for image-text linkages present in English. This may result in unexpected behavior when evaluating sarcasm in multi-lingual or code-mixed settings. Thus, we advocate the safe use of these approaches with human-in-the-loop validation. Similarly, we release all code and prompt templates to support transparency, reproducibility, and further research on coherence-aware reasoning in multi-modal models.
Acknowledgments
This research was in part supported by the U.S. National Science Foundation (NSF) award #1820609. Part of the research results were obtained using the computational resources provided by CloudBank (https://www.cloudbank.org/), which was supported by the NSF award #1925001.
Appendix
Appendix A Distribution of Predicted Coherence Relations
As an extended analysis, we present the distribution of predicted Coherence Relations by GPT4o coupled with IRONIC in Tables 3 and 4. This distribution has been constructed by parsing the Relation Extraction reasoning step. In MMSD 2.0 and RedEval, eight and two image-text pairs respectively do not have predicted Coherence Relations because their rationales exceeded the maximum token limit.
CR Non-Sarcastic Sarcastic
Insertion 122 75
Extension 248 297
Concretization 84 39
Projection 458 283
Restatement 456 339
Table 3: Distribution of CRs in the MMSD 2.0 Dataset
CR Non-Sarcastic Sarcastic
Insertion 51 24
Extension 96 129
Concretization 27 16
Projection 101 75
Restatement 334 149
Table 4: Distribution of CRs in the RedEval Dataset
Appendix B Model Availability
This section focuses on the details regarding model availability and parameters used for evaluation. For all models, we set temperature to 00 or do_sample=False, maximum output tokens to $512$ and the random seed set to $42$, wherever possible to ensure reproducibility.
B.1 Proprietary Models
OpenAI GPT:
We access the GPT-4o model via a custom deployment using Azure OpenAI. We evaluate gpt-4o-2024-11-20 with a custom safety filter to restrict content of high severity level. We also use the public version of GPT-4o, via the official OpenAI API.
B.2 Open-Source Models
We evaluate InternVL3 14B MLLM (OpenGVLab/InternVL3-14B on Huggingface) using the VLLM  [^1] https://github.com/vllm-project/vllm framework.
Appendix C Prompt Templates
We use various prompt templates with different system/user messages for IRONIC and the baselines mentioned in Section 4.3. The different prompts and system messages used are present in the appendix.
Appendix D Postprocessing MLLM Responses
Since both datasets feature image-caption pairs on a wide selection of topics, around 50+ images were flagged by Azure GPT4o’s strict safety filter or rejected by the model during evaluation. To ensure test set consistency, we decided to rerun these samples using the official OpenAI API instead and were able to get results.
System Message for Baseline Zero-Shot/CoT Evaluation You are an expert linguist. Your task is to analyze a image-text pair and determine if it is ’sarcastic’ or ’non-sarcastic’.
System Message for IRONIC You are an expert linguist, and your task is to analyze and predict the most applicable Coherence Relations for image-text pairs. A coherence relation describes the structural, logical, and purposeful relationships between an image and its caption, capturing the author’s intent. These are the possible coherence relations you can assign to an image-text pair: - Insertion: The salient object described in the image is not explicitly mentioned in the text. - Concretization: Both the text and image contain a mention of the main visual entity. - Projection: The main entity mentioned in the text is implicitly related to the visual objects present in the image. - Restatement: The text directly describes the image contents. - Extension: The image expands upon the story or idea in the text, presenting new elements or elaborations, effectively filling in narrative gaps left by the text.
Baseline Zero-Shot Prompt System <insert-system-message> User Output 0 if it does not contain sarcastic content, and 1 if it does. Do not add anything else in your response. <insert-image-text-pair>
Baseline Zero-Shot CoT Prompt System <insert-system-message> User Let’s think step-by-step and analyze the relationship between the text and image carefully. <insert-image-text-pair> Assistant Analysis: <add-analysis-from-model> User Using your rationale, please determine if the image-text pair is sarcastic in nature. Output 0 if it does not contain sarcastic content, and 1 if it does. Do not add anything else in your response.
$S^{3}$ Agent Prompt User Given the following image and text, please judge whether there is sarcasm based on the 3 perspectives below. 1. Superficial Expression: This includes detect underlying critiques in contexts through image-text discrepancies. 2. Semantic Information: This includes detect extreme portrayals and metaphorical in contexts through image-text semantic. 3. Sentiment Expression: This includes detect criticize emotion on specific subjects or behaviors in the content. Without considering conclusions drawn solely from images or text, both must be considered together. Then, you should output the corresponding chain-of-thought to support your answer. <insert-image-text-pair> Assistant Analysis: <add-analysis-from-model> User Unmask the hidden intent! Given the above rationales, delve into its layers of meaning. Analyze the surface - the literal words used. Pierce deeper to uncover the semantic information - the intended meaning behind those words. Finally, gauge the sentiment - the emotional undercurrent. By weaving these insights together, can you crack the code of sarcasm and determine if the comment is meant to be sincere or laced with sarcastic? Follow these rules: 1. If any perspective cannot determine sarcasm due to lack of information, disregard that perspective. 2. If any of the views conflict, choose the view with the most well-founded reasoning. 3. Output 0 if it does not contain sarcastic content, and 1 if it does. Do not add anything else in your response.
IRONIC Prompt System <insert-system-message> User Let’s think step-by-step and analyze the relationship between the text and image carefully. Please also identify the most appropriate coherence relation between the text and image. <insert-image-text-pair> Assistant Analysis: <add-analysis-from-model> User Using the coherence relation analysis and the input image-text pair, please determine if the content contains multi-modal sarcasm. Content where both image and text together create a sarcastic effect is considered multi-modal sarcasm, do not focus on the sarcasm in only the image or text. Clearly distinguish simple humor and satire which is not sarcastic from mockery and irony which is sarcastic. Output 0 if it does not contain multi-modal sarcasm, and 1 if it does. Do not add anything else in your response.